# Generative Artificial Intelligence for Beamforming in Low-Altitude Economy

### v2

@@ III-A4 Transformers
Transformers leverage global context through self-attention mechanisms, which generate content by modeling the conditional probability of each element in the sequence. A key advantage of self-attention lies in its ability to directly capture long-range dependencies and relationships within sequential data, rather than relying on fixed-size windows or recurrent connections. Consequently, the inherent capability for modeling comprehensive spatiotemporal relationships makes Transformers particularly valuable for tasks requiring strong temporal and spatial awareness. For example, in [14], the authors embedded dynamic network states into structured tokens and applied a Transformer-based model to model multi-scale spatiotemporal dependencies. The resulting representations effectively captured time-varying channel dynamics and user preferences, thereby enhancing the exploration capability of the actor-critic network in deep reinforcement learning (DRL). In addition, the robustness of Transformers derives from the self-attention mechanism and their capacity to process non-sequential, variable-length, and partially corrupted input. Specifically, self-attention enables the model to dynamically reweight input tokens based on contextual relevance, thus allowing suppression of noisy features and inference of missing information by referencing global context. Consequently, the robustness of Transformers makes them particularly well suited for handling irregular input streams, fluctuating channel quality, and asynchronous data transmissions in LAE scenarios.
III-B Generative AI for Beamforming in Low-Altitude Economy
GenAI elevates beamforming in LAE by dynamically adapting to complex environmental uncertainties and collaborative network demands. This section demonstrates its capabilities in mission-critical tasks and resilient performance enhancement.
@@ III-B1 Communication Tasks
The mobility and interference of LAE demand agile beamforming. GenAI transcends conventional approaches, thus enabling robust task execution, as detailed in two key tasks.
  • Channel Modeling and Estimation: Compared to conventional communication scenarios, LAE places higher demands on channel modeling and estimation due to its complex and highly dynamic environment. Specifically, factors such as adverse weather conditions, obstruction by buildings or terrain features, interference from surrounding nodes, and the rapid mobility of LAPs significantly impact signal quality. These factors often lead to signal attenuation and pronounced multipath propagation effects, thus posing substantial challenges to accurate channel modeling and estimation. Although conventional beamforming techniques can adapt to certain channel variations, they struggle to effectively address the rapid temporal and spatial fluctuations introduced by these environmental factors. Therefore, there is a clear need for innovative approaches, and GenAI offers promising solutions by leveraging advanced data-driven frameworks to improve channel modeling and estimation.
    For instance, Zhang et al. [11] developed a GAN-based model to improve channel estimation in millimeter-wave communications for low-altitude unmanned aerial vehicle (UAV) networks. Specifically, the model allows each UAV to independently train channel models for beamforming directions, with samples shared through a collaborative framework to build a generalized model, which improves estimation accuracy. Additionally, Rasheed et al. [12] proposed a distributed GAN model based on long short-term networks for low-altitude maritime UAV communications. Specifically, UAVs utilize the model to extract real-time CSI and generate synthetic samples, which improve the accuracy of channel state estimation.
@@
  • Resource Allocation: Compared to conventional communication environments, efficient spectrum and power resource allocation in LAE scenarios is significantly more challenging due to dense node deployments, rapid network topology changes caused by node mobility, and intense competition for limited resources. However, conventional beamforming approaches often fail to quickly adjust beam configurations in response to these complex factors, thus leading to inefficient resource management. In contrast, GenAI-based beamforming methods explicitly link beamforming optimization with resource allocation, enabling simultaneous and real-time adjustment of beam directions and spectrum assignments to address resource conflicts and enhance utilization efficiency.
    For instance, a VAE-based framework can encode network topology variations and resource competition conditions into concise latent features. Subsequently, the decoder directly generates optimized beamforming parameters and adaptive spectrum allocation schemes. This integrated approach effectively mitigates interference, resolves resource competition conflicts, and improves overall spectrum and power utilization efficiency in highly dynamic LAE scenarios.
@@
III-B2 Performance Metrics
LAE demands adaptive security and energy resilience against evolving threats. GenAI delivers fundamental breakthroughs, as detailed in the following two aspects.
  • Communication Security: Compared to the conventional communication scenarios, communication security in LAE is particularly critical. On the one hand, different tasks may rely on different networks or frequency bands, which means that the LAPs frequently need to switch between multiple communication networks and frequency bands. As a result, this increases the vulnerability of communication links to network attacks or malicious interference. On the other hand, LAE includes tasks in critical areas such as logistics and emergency rescue, where the reliability of communication security directly impacts the efficiency and safety of these services. However, conventional beamforming algorithms rely solely on signal strength to optimize beam direction, which limits their ability to handle dynamic malicious interference. In contrast, GenAI models excel at analyzing latent data distributions and dynamically optimizing anti-jamming strategies under uncertain conditions.
    For example, in [10], the authors combined a diffusion-based optimization framework with a mixture of expert (MoE)-Transformer network, and leveraged multi-head attention to model spatiotemporal channel dependencies. Moreover, this method dynamically optimizes the three-dimensional beamforming to resist eavesdropping interference, mitigate location errors, and address channel uncertainties, thus enhancing the anti-jamming resilience and secrecy performance in the mission-critical low-altitude communication systems.
  • Energy Consumption: In LAE senarios, the frequent beam redirection by LAPs, interference suppression between high-density nodes, and multi-network handovers significantly increase LAP energy consumption compared to conventional communication environments. However, conventional beamforming algorithms rely on pre-set power distribution strategies, which hinder the ability to balance communication quality and energy efficiency in dynamic topology and channel fluctuation scenarios. GenAI can learn the relationship between environmental features and energy consumption online, autonomously generating beam control strategies that optimize energy efficiency and overcome the limitations of traditional static approaches.
    For instance, a beam control system based on GAN can employ a generator to simulate low-power beam patterns and dynamically optimize beam width and radiation power. Meanwhile, the discriminator can assess signal integrity based on real-time channel states, which ensures a reduction in energy consumption without sacrificing performance.
@@ III-C Lessons Learned
  2. 2.
    GenAI significantly improves channel modeling and estimation in LAE environments. Specifically, GenAI models excel at learning intricate channel characteristics even with limited data, thus enabling accurate real-time CSI extraction. In addition, GenAI models also fundamentally redefine resource allocation. In particular, GenAI-enhanced beamforming methods can simultaneously adjust signal directions and spectrum assignments. Consequently, the capability of GenAI ensures efficient resource management and effectively mitigates interference in highly dynamic conditions of LAE.
  3. 3.
@@
IV Case Study: Generative AI for aerial collaborative beamforming in Low-Altitude Economy
@@ IV-D Performance Evaluation
To validate the effectiveness of the proposed GDMTD3 algorithm, extensive simulations were conducted and compared with baseline algorithms, including deep deterministic policy gradient (DDPG), soft actor-critic (SAC), and standard TD3. The experimental findings are summarized as follows.
As shown in Fig. 5(a), the GDMTD3 algorithm achieved higher and more stable rewards across 1000 training iterations compared to the baseline methods. This demonstrates its ability to learn optimal policies effectively and adapt to dynamic environments, while the baseline algorithms exhibited slower convergence and greater variability.
In terms of optimization objectives, Fig. 5(b) shows that the proposed GDMTD3 achieved the highest secrecy rate with the lowest energy consumption per step, effectively balancing the security and efficiency. In contrast, SAC and TD3 showed moderate performance in one metric but failed to achieve the optimal trade-off.

### v3

@@ III-A4 Transformers
Transformers leverage global context through self-attention mechanisms, which generate content by modeling the conditional probability of each element in the sequence. A key advantage of self-attention lies in its ability to directly capture long-range dependencies and relationships within sequential data, rather than relying on fixed-size windows or recurrent connections. Consequently, the inherent capability for modeling comprehensive spatiotemporal relationships makes Transformers particularly valuable for tasks requiring strong temporal and spatial awareness. For example, in [14], the authors embedded dynamic network states into structured tokens and applied a Transformer-based model to model multi-scale spatiotemporal dependencies. The resulting representations effectively captured time-varying channel dynamics and user preferences, thereby enhancing the exploration capability of the actor-critic network in deep reinforcement learning (DRL). In addition, the robustness of Transformers derives from the self-attention mechanism and their capacity to process non-sequential, variable-length, and partially corrupted input. Specifically, self-attention enables the model to dynamically reweight input tokens based on contextual relevance, thus allowing the model to suppress noisy features and infer missing information by referencing global context. Consequently, the robustness of Transformers makes them particularly well suited for handling irregular input streams, fluctuating channel quality, and asynchronous data transmissions in LAE scenarios.
III-B GenAI for Beamforming in Low-Altitude Economy
GenAI offers a holistic enhancement to beamforming in the LAE, fundamentally improving both core communication tasks and overall system performance.
@@ III-B1 Communication Tasks
The dynamic nature of LAE communication demands highly agile and intelligent management. To this end, GenAI provides a powerful framework for robustly executing critical communication tasks.
  • Channel Estimation: Compared to conventional communication scenarios, there is an increased demand for channel estimation of LAE due to the complex and highly dynamic environment. Specifically, some factors such as adverse weather conditions, obstruction by buildings and the rapid mobility of LAPs often lead to signal attenuation and pronounced multipath propagation effects, thus posing substantial challenges to accurate channel estimation. In this case, GenAI models can offer a full-cycle solution by addressing different stages of the channel information processing. The initial stage involves incomplete historical channel information reconstruction, where GANs can be used to generate realistic samples to fill information gaps. Moreover, for the current channel state analysis, VAEs can provide a probabilistic assessment by modeling the channel distribution. Furthermore, GDMs can refine noisy observations through iterative reverse diffusion. For instance, in [10], the authors proposed a GDM-based model that frames channel estimation as a denoising problem and recovers the true channel by reversing a noise-adding process. In addition, for future channel information prediction, Transformers can leverage self-attention to capture temporal correlations within sequential data. For example, Ju et al. [13] developed a Transformer-based framework that uses a sequence of previous channel states to predict future channel parameters and thereby mitigate channel aging.
@@
  • Resource Allocation: Compared to conventional communication environments, efficient spectrum and power resource allocation in LAE scenarios is more challenging due to dense node deployments, rapid network topology changes caused by node mobility, and intense competition for limited resources. However, conventional beamforming approaches often fail to quickly adjust beam configurations in response to these complex factors, thus leading to inefficient resource management. In contrast, GenAI-based beamforming methods explicitly link beamforming optimization with resource allocation, enabling simultaneous and real-time adjustment of beam directions and spectrum assignments to address resource conflicts and enhance utilization efficiency. For instance, a VAE-based framework can encode network topology variations and resource competition conditions into concise latent features. Subsequently, the decoder directly generates optimized beamforming parameters and adaptive spectrum allocation schemes. This integrated approach effectively mitigates interference, resolves resource competition conflicts, and improves overall spectrum and power utilization efficiency in highly dynamic LAE scenarios.
III-B2 Performance Metrics
LAE demands adaptive security and energy resilience against evolving threats. To address this demand, GenAI shows promising progress, which can be described as follows.
  • Communication Security: Compared to the conventional communication scenarios, communication security in LAE is particularly critical. On the one hand, different tasks may rely on different networks or frequency bands, which means that the LAPs frequently need to switch between multiple communication networks and frequency bands. As a result, this increases the vulnerability of communication links to network attacks or malicious interference. On the other hand, LAE includes tasks in critical areas such as logistics and emergency rescue, where the reliability of communication security directly impacts the efficiency and safety of these services. However, conventional beamforming algorithms rely solely on signal strength to optimize beam direction, which limits their ability to handle dynamic malicious interference. In contrast, GenAI models excel at analyzing latent data distributions and dynamically optimizing anti-jamming strategies under uncertain conditions. For example, in [11], the authors combined a diffusion-based optimization framework with a mixture of expert (MoE)-Transformer network, and leveraged multi-head attention to model spatiotemporal channel dependencies. Moreover, this method dynamically optimizes the three-dimensional beamforming to resist eavesdropping interference, mitigate location errors, and address channel uncertainties, thus enhancing the anti-jamming resilience and secrecy performance in the mission-critical low-altitude communication systems.
  • Energy Consumption: In LAE scenarios, the frequent beam redirection by LAPs, interference suppression between high-density nodes, and multi-network handovers significantly increase LAP energy consumption compared to conventional communication environments. However, conventional beamforming algorithms rely on preset power distribution strategies, which hinder the ability to balance communication quality and energy efficiency in dynamic topology and channel fluctuation scenarios. GenAI can learn the relationship between environmental features and energy consumption online, autonomously generating beam control strategies that optimize energy efficiency and overcome the limitations of traditional static approaches. For instance, a beam control system based on GAN can employ a generator to simulate low-power beam patterns and dynamically optimize beam width and radiation power. Meanwhile, the discriminator can assess signal integrity based on real-time channel states, which ensures a reduction in energy consumption without sacrificing performance.
@@ III-C Lessons Learned
  • GenAI significantly improves channel estimation in LAE environments. Specifically, GenAI models can offer a full-cycle solution that addresses different stages of the estimation process. In addition, GenAI models also fundamentally redefine resource allocation. In particular, GenAI-enhanced beamforming methods can simultaneously adjust signal directions and spectrum assignments.
@@
IV Case Study: GenAI for collaborative beamforming in Low-Altitude Economy
@@ IV-D Performance Evaluation
To validate the effectiveness of the proposed GDMTD3 algorithm, extensive simulations were conducted and compared with baseline algorithms, including deep deterministic policy gradient (DDPG), soft actor-critic (SAC), and standard TD3.
As shown in Fig. 5(a), the GDMTD3 algorithm achieved higher and more stable rewards across 1000 training iterations compared to the baseline methods. This demonstrates its ability to learn optimal policies effectively and adapt to dynamic environments, while the baseline algorithms exhibited slower convergence and greater variability. In terms of optimization objectives, Fig. 5(b) shows that the proposed GDMTD3 achieved the highest secrecy rate with the lowest energy consumption per step, effectively balancing the security and efficiency. In contrast, SAC and TD3 showed moderate performance in one metric but failed to achieve the optimal trade-off.